OpenStudy (anonymous):

A die with 6 sides is rolled. What is the probability of rolling a number less than 5?

OpenStudy (shubhamsrg):

whats your sample space and favorable outcomes ?

OpenStudy (anonymous):

My options are 2/3, 5/6, 4, or 1/6

OpenStudy (e.cociuba):

I'm gunna go w/ 2/3 ;))

OpenStudy (anonymous):

so it cant be 6 or 5.

OpenStudy (anonymous):

0<5

OpenStudy (e.cociuba):

Number less than 5 will be 1,2,3,4 n(s)=6 p(s)=4 p(E)= 4/6=2/3

OpenStudy (johnweldon1993):

Good job @e.cociuba ;P

OpenStudy (e.cociuba):

Mmhmm I couldn't say the same for u @johnweldon1993 lol ;P

OpenStudy (anonymous):

so four options, 1, 2, 3, 4 so 4/6 = 2/3 so there is a 66% chance of rolling a number 1 - 4. so a bookmaker would try and make u bet at a price less than $1.51, or offer you 3-2 odds on, but only a fool would bet becuase it would be under the odds and u would lose long term. Never gamble at a casino, you will lose.

OpenStudy (anonymous):

Thanks for the help! I'm taking statistics and I'm terrible at it. I have a few more questions.

OpenStudy (e.cociuba):

Np! :) Post them separatly though. It's easier tht way :D

OpenStudy (anonymous):

A spinner has equal regions numbered 1 through 21. What is the probability that the spinner will stop on an even number or a multiple of 3?

OpenStudy (anonymous):

My options are 10/9, 2/3, 1/3, or 17

OpenStudy (e.cociuba):

The even numbers or a multiple of three are: 2,3,4,6,8,9,10,12,14,15,16,18,20,21 14/21 = 2/3

OpenStudy (anonymous):

Which of the following cannot be a probability? 1/2, 1, -1, or 0